# Data-driven modeling of multiscale phenomena with applications to fluid turbulence

### v1

@@ Abstract
A top-down effective theory can be derived formally from the corresponding fundamental theory only in very specific limits, e.g., the Chapman-Enskog theory requires the gas to be near local thermodynamic equilibrium and the Knudsen number (characterizing the mean free path) to be small. However, one can ask whether a sufficiently general approach exists that, at least in some classes of problems, allows an effective theory to be inferred in a systematic way from the corresponding fundamental theory. A particular class of problems where this capability would be most welcome involves multiscale phenomena, such as turbulence in fluids and plasmas, which can feature structures with length and time scales varying by tens of orders of magnitudes.
@@
To circumvent this limitation, various models of fluid turbulence have been introduced which involve either temporal or spatial coarse-graining as in, respectively, the Reynolds-averaged Navier-Stokes (RANS) or large eddy simulation (LES). While the coarse-graining procedure itself is formal and general, the equations describing the averaged fields and their fluctuations cannot be closed without making specific assumptions, yielding a phenomenological description. In the case of both RANS and LES, an additional closure term appears in the momentum equation. It represents the effect of the small (modeled) scales on the large (resolved) ones and its functional form is assumed based on various empirical arguments rather than derived.
Figure 1: Representative flow fields used to generate the test data. Shown is the vorticity field.
@@
This letter introduces a general data-driven framework for inferring an effective field theory of multiscale phenomena from the underlying fundamental description. This framework requires no specific phenomenological assumptions and ensures that the resulting theory is equivariant and interpretable. Unlike the bulk of SGS models, the dynamics of small scales and scale interaction are characterized explicitly, which enables an accurate description of backscatter in the case of fluid turbulence.
Methodology—We will retain the coarse-graining approach used in LES to separate the degrees of freedom with length scales larger and smaller than an arbitrary cutoff scale $\Delta$. The choice of the kernel $G$ in equation (2) is partly arbitrary but must satisfy a number of constraints such as locality in physical and Fourier space [19, 20]. We use a normalized Gaussian kernel (with second moment $\sigma^{2}=\Delta^{2}/12$) exclusively in this letter as it maximizes locality in both spaces. To make the description finite-dimensional (e.g., to enable numerical solution), we additionally apply a sharp spectral filter, which cuts off scales past $\Delta$, making the combined filter non-invertible. The resulting splitting formally defines the governing equations for both large scales (here the system (3)) and small scales. The variables describing small scales and the associated governing equations are discarded in LES—which leads to the problems discussed previously—but are occasionally retained in RANS. In the latter case, the effect of small scales is represented by an SGS stress tensor whose formal evolution equation [21] includes variables other than $\bar{u}_{i}$, $\bar{p}$ and $\tau_{ij}$, generated by the nonlinearity, so that the resulting system of equations is not closed.
@@
of the SGS tensor into the Leonard ($L$), Reynolds ($R$), and cross ($C$) stress tensors describing, respectively, the interaction of large scales, interaction of small scales, and interscale interaction. While both $\bf u$} and $\bf u$$^{\prime}$ (and thereby $L$, $C$ and $R$) can be reconstructed from $\bar{\bf u}$ exactly for the Gaussian filter, the sharp spectral filter is not invertible and so some of the tensors may not be accurately described in terms of the resolved fields.
Both the governing equations for the small-scale variables and the constitutive relation for the SGS stress tensor can be inferred using a data-driven approach. Specifically, we use the SPIDER framework [24] which has been thoroughly validated by recovering fundamental models describing active nematics [25], turbulent channel flow [26], and turbulent convection [27]. SPIDER combines group representation theory, weak formulation of differential equations, and sparse regression to infer complete equivariant models, including boundary conditions, directly from data (experimental or numerical).
SPIDER searches for sparse relations explicitly equivariant with respect to a relevant symmetry group. This is done systematically and comprehensively by exploring search spaces corresponding to different irreducible representations. For instance, for the rotation group, each search space is defined by a basis set of tensors $\{F^{1}_{k},F^{2}_{k},\cdots\}$ that transform according to an irreducible representation $k$ and are constructed using contractions and tensor products of the variables (e.g., $\overline{u}_{i},\overline{p},\cdots$) and their derivatives (e.g., $\partial_{t}\overline{u}_{i},\nabla_{i}\overline{p},\nabla_{j}\overline{u}_{i},\cdots$). The tensor libraries $\mathcal{L}_{k}=\{F^{1}_{k},\cdots,F^{N_{k}}_{k}\}$ are then finitely truncated by selecting an upper bound on the complexity of terms, typically defined as the number of variables and differential operators appearing in a given term. In this way, multiple physical relations can be identified in the form
$c_{1}F^{1}+\cdots+c_{N}F^{N}=0,$ (5)
where $c_{n}$ are learned constant coefficients.
In this letter, we consider non-relativistic flows, so the relevant symmetries include translations, rotations, and Galilean transformations. For simplicity, we consider only the integer-rank tensor (not spinor) representations of the rotation group. Evaluating equation (5) in weak form over a collection of spatiotemporal regions yields an overdetermined linear system $Q{\bf c}=0$ which is solved to find an approximate sparse solution ${\bf c}=(c_{1},\cdots,c_{N}$). Depending on the objective, one can either require $|{\bf c}|=1$ (homogeneous regression) or $c_{m}=1$ for some $1\leq m\leq N$ (inhomogeneous regression). Finally, the scaling of the coefficients with parameters (here, $\Delta$ and $\nu$) is inferred through dimensional analysis and verified by generating data for a range of parameters.
@@
We consider two-dimensional turbulence, which features pronounced coherent structures (eddies) that break statistical isotropy/homogeneity as well as strong backscatter [28], making it an ideal testing ground. We use a pseudospectral solver [11] to generate fully resolved DNS data on a square domain of size $\ell=2\pi$ with periodic boundary conditions. Figure 1 shows snapshots of such flows representing the inverse cascade (F1), freely decaying turbulence (F2) and the direct cascade (F3) used as initial conditions. For each of the three initial conditions, a numerical solution corresponding to $O(1)$ eddy turnover time is generated for each value of the viscosity corresponding to Reynolds number $Re=u_{rms}\ell/\nu$ equal to $10^{4}$, $10^{5}$, and $10^{6}$ on computational grids with respective resolutions of $2048^{2}$, $4096^{2}$, and $4096^{2}$. Next, coarse-grained datasets are generated by applying the filtering operator for a range of cutoff scales $\Delta$. Finally, equations are inferred from the coarse-grained data using the Python implementation of SPIDER [24].
@@
Results—To determine whether the SGS stress tensor can be represented by functions of the resolved fields $\bar{\bf u}$ and $\bar{p}$, we use them to construct the respective rank-0 and symmetric trace-free rank-2 libraries and employ inhomogeneous regression. Across all data sets we consistently find
$\tau_{ij}$ $\approx c_{1}\nabla_{k}\bar{u}_{i}\nabla_{k}\bar{u}_{j}+c_{2}\nabla_{k}\nabla_{m}\bar{u}_{i}\nabla_{k}\nabla_{m}\bar{u}_{j},$ (6)
where the second right-hand-side term is generally small compared to the first one. Incidentally, this result corresponds to the formal series expansions of $L+C$ in powers of $\Delta$ [20]. Using dimensional analysis and the numerical values of the coefficients identified by SPIDER, we find $c_{1}=\Delta^{2}/12$ and $c_{2}=\Delta^{4}/288$, consistent with the coefficients of the formal expansion.
@@
Figure 2: Representative snapshot of the energy flux $\Pi$ for DNS (a), the LCR model (b), DS model (c), and DM model (d) for the flow F2 at $Re\sim 10^{6}$ and cutoff scale $\Delta=\ell/64$.
On the other hand, the tensor $R$, which is generally small compared with $L$ and $C$, cannot be expressed in terms of $\bar{\bf u}$ or $\bar{p}$ with even moderate accuracy. Therefore, we simply treat $R$ as an additional variable in the effective field theory, one that describes small scales, and parameterize the SGS tensor as
$\tau_{ij}\approx\frac{\Delta^{2}}{12}\nabla_{k}\bar{u}_{i}\nabla_{k}\bar{u}_{j}+\frac{\Delta^{4}}{288}\nabla_{k}\nabla_{m}\bar{u}_{i}\nabla_{k}\nabla_{m}\bar{u}_{j}+R_{ij}$ (7)
@@
which we refer to as the LCR model. This model not only reproduces the individual components of the SGS stress tensor faithfully but also accurately captures the spatiotemporal structure of the energy flux $\Pi\equiv-\tau_{ij}\,\bar{S}_{ij}$ (where $S_{ij}$ denotes the symmetric part of $\nabla_{i}u_{j}$) including backscatter (regions where $\Pi<0$). This is illustrated in figure 2 for the freely decaying flow F2. For reference, the figure also shows the predictions of the widely used dynamic Smagorinsky (DS) model as well as the dynamic mixed (DM) model (see End Matter), which is considered to be one of the most accurate LES models [29]. Both phenomenological models fail to predict local fluxes.
@@
The inclusion of $R$ as a new field requires an additional governing equation in order to close the system. This equation can be inferred using rank-0 and symmetric trace-free rank-2 libraries constructed from $\bar{u}_{i}$, $\bar{p}$, and $R_{ij}$. We robustly discover an evolution equation of the form
$\partial_{t}R_{ij}$ $=c_{1}\bar{u}_{k}\nabla_{k}R_{ij}+c_{2}(R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i})$
@@
Using a combination of dimensional analysis and numerical values of the coefficients for different data sets, we find $c_{1}=-1$, $c_{2}=1$ and $c_{3}=\nu$. For $c_{4}$, we find that it has both dimensions and numerical values consistent with the inverse of the eddy turnover time, exhibiting no clear dependence on the parameters $\nu$ and $\Delta$. Therefore, we relax the assumption of translational invariance made by SPIDER that requires $c_{4}$ to be independent of position and time and take it to be a function of the resolved flow. For instance, in the inviscid limit, rescaling the velocity of the initial condition by a constant factor changes the inverse time scale and, hence, $c_{4}$ by the same factor, so $c_{4}$ should be linear in $\bar{u}$. Because $c_{4}$ is a scalar, it should be expressible as a function of the energy $E=u_{i}u_{i}/2$ as well as the invariants of the tensor $\nabla_{i}\bar{u}_{j}$ [30]; in 2D, these are given by $|\bar{S}|=\sqrt{2\bar{S}_{ij}\bar{S}_{ij}}$ and $|\bar{\Omega}|=\sqrt{2\bar{\Omega}_{ij}\bar{\Omega}_{ij}}$, where ${\Omega}_{ij}$ is the antisymmetric component of $\nabla_{i}u_{j}$. Dimensional arguments would then require $c_{4}=\Delta^{-1}\sqrt{E}f(\Delta|\bar{\Omega}|/\sqrt{E},\Delta|\bar{S}|/\sqrt{E},\cdots)$, where $f$ is an arbitrary function. A general relation of this form can be inferred from the data using, e.g., symbolic regression. However, a good approximation can be found using a series expansion of $f$:
@@
$\partial_{t}R_{ij}$ $+\bar{u}_{k}\nabla_{k}R_{ij}=R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i}$ (10)
@@
To measure the accuracy in predicting a tensor variable, we will employ a magnitude-aware correlation
between tensor fields $A$ (ground truth) and $B$ (prediction), where $\langle\cdot\rangle$ denotes spatial averaging, with $\mathcal{C}(A,B)=1$ only when $A=B$. Using $C_{R}=\mathcal{C}(\partial_{t}R_{\text{DNS}},\partial_{t}R_{\text{LES}})$ as a metric, we find equation (10) to accurately describe all flow regimes for sufficiently small $\Delta$ (cf. figure 3). As expected, the accuracy deteriorates as $\Delta$ approaches the scale of coherent structures, since the coarse-grained description is not meant to model these.
Figure 3: The accuracy of the evolution equation (10), quantified by the correlation $C_{R}$, as a function of $\Delta/\ell$ for F1 (blue), F2 (orange) and F3 (yellow) at $Re=10^{6}$.
Conclusions—Taken together, the closed system of equations (3), (7) and (10) defines an effective field theory which describes incompressible fluid turbulence up to the cutoff scale $\Delta$. By construction, this theory is explicitly equivariant with respect to translations and rotations but also turns out to be explicitly Galilean-invariant, as it should. Not only is the resulting description more accurate and general than state-of-the-art LES models, it is explicit and easily interpretable, which allows one to make specific predictions. For example, at scales $\Delta$ small compared to the size of large-scale coherent structures, the energy flux is dominated by the contribution from the $C$ tensor and scales as $\Delta^{4}$, which is not directly implied by the fundamental model (1).
It is worth emphasizing that the data-driven framework described in this letter requires a minimal number of ingredients—the fundamental model (which could also be derived in a data-driven way [26]), some of the fundamental symmetries of the problem, and dimensional analysis—and no specific phenomenological assumptions. Hence, it should readily generalize to other multiscale problems with polynomial nonlinearities.

### v2

@@ Abstract
A top-down effective theory can be derived formally from the corresponding fundamental theory only in very specific limits, e.g., the Chapman-Enskog theory requires the gas to be near local thermodynamic equilibrium and the Knudsen number (characterizing the mean free path) to be small. However, when formal derivations fail, one can ask whether a sufficiently general approach exists that, at least in some classes of problems, allows an effective theory to be inferred in a systematic way from the corresponding fundamental theory. A particular class of problems where this capability would be most welcome involves multiscale phenomena, such as turbulence in fluids and plasmas, which can feature structures with length and time scales varying by tens of orders of magnitudes.
@@
To circumvent this limitation, various models of fluid turbulence have been introduced which involve either temporal or spatial coarse-graining as in the Reynolds-averaged Navier-Stokes (RANS) or large eddy simulation (LES), respectively. While the coarse-graining procedure itself is formal and general, the equations describing the averaged fields and their fluctuations cannot be closed without making specific assumptions, yielding a phenomenological description. In the case of both RANS and LES, an additional closure term appears in the momentum equation. It represents the effect of the small (modeled) scales on the large (resolved) ones and its functional form is assumed based on various empirical arguments rather than derived.
@@
Figure 1: Representative flow fields used to generate the test data. Shown is the initial vorticity field $\omega$ .
@@
This letter introduces a general data-driven framework for inferring an effective field theory of multiscale phenomena from the underlying fundamental description. This framework requires no phenomenological assumptions and ensures that the resulting theory is equivariant and interpretable. To illustrate this framework, we apply it to fluid turbulence and infer a subgrid-scale model that accurately describes backscatter. This is due, in no small part, to the model’s explicit characterization of the subgrid scales.
Methodology—We will retain the coarse-graining approach used in LES to separate the degrees of freedom with length scales larger and smaller than a cutoff scale $\Delta$; while $\Delta$ is an adjustable parameter of the model, it must be at a scale substantially smaller than the scale of coherent structures, as SGS models are not designed to capture these. The choice of the kernel $G$ in equation (2) is partly arbitrary but must satisfy a number of constraints such as locality in physical and Fourier space [19, 20]. We use a normalized Gaussian kernel (with second moment $\sigma^{2}=\Delta^{2}/12$) exclusively in this letter as it maximizes locality in both spaces. To make the description finite-dimensional (e.g., to enable numerical solution), we additionally apply a sharp spectral filter, which cuts off scales past $\Delta$, making the combined filter non-invertible. The resulting splitting formally defines the governing equations for both large scales (here the system (3)) and small scales. The variables describing small scales and their associated governing equations are discarded in LES—which leads to the problems discussed previously—but are occasionally retained in RANS. In the latter case, the effect of small scales is represented by an SGS stress tensor whose formal evolution equation [21] includes variables other than $\bar{u}_{i}$, $\bar{p}$ and $\tau_{ij}$, generated by the nonlinearity, so that the resulting system of equations is not closed.
@@
of the SGS tensor into the Leonard ($L$), Reynolds ($R$), and cross ($C$) stress tensors describing, respectively, the interaction of large scales, interaction of small scales, and interscale interaction. While both $\bf u$} and $\bf u$$^{\prime}$ (and thereby $L$, $C$ and $R$) can be reconstructed from $\bar{\bf u}$ exactly for the Gaussian filter, the sharp spectral filter is not invertible and so some of the tensors may not be accurately reconstructed in terms of the resolved fields.
On a sufficiently fine grid and for kernels with finite moments, any filtered field can be approximated using the moment expansion [20] as a series in $\Delta/\ell_{c}$, where $\ell_{c}$ is a characteristic length scale such as a vortex size. In particular, one finds $L=O((\Delta/\ell_{c})^{2})$, $C=O((\Delta/\ell_{c})^{4})$ and $R=O((\Delta/\ell_{c})^{6})$, which means that $L$ provides the dominant contribution to $\tau$ while $C$ and $R$ yield progressively smaller corrections. The leading order term of the moment expansion representation of $\tau$ (and $L$) yields the nonlinear gradient model [32] that was recently rediscovered using a data-driven approach [14]. While this model yields a reasonably accurate approximation of the SGS tensor, it incorrectly predicts that the energy flux $\Pi\equiv-\tau_{ij}\,\bar{S}_{ij}$ (where $S_{ij}$ denotes the symmetric part of $\nabla_{i}u_{j}$) vanishes identically. A correct description of the fluxes requires a far more accurate approximation of the SGS tensor.
@@
We consider two-dimensional turbulence, which features pronounced coherent structures (eddies) that break statistical isotropy/homogeneity and is characterized by strong backscatter [28], making it an ideal testing ground. We use a pseudospectral solver jax-cfd [34] to generate fully resolved DNS data on a square domain of size $\ell=2\pi$ with periodic boundary conditions. Figure 1 shows snapshots of such flows representing the inverse cascade (F1), freely decaying turbulence (F2) and the direct cascade (F3) used as initial conditions. For each of the three initial conditions, a numerical solution corresponding to about one eddy turnover time is generated on a computational grid with resolution of $2048^{2}$, $4096^{2}$, and $4096^{2}$ for $\nu=10^{-4}$, $10^{-5}$, and $10^{-6}$, respectively. The corresponding Reynolds numbers $Re=u_{\rm RMS}\ell/\nu$ are reported in table 1. The initial conditions and the software for generating the data are available online [36].
Next, the numerical solutions are coarse-grained by applying the filtering operator over a range of cutoff scales $\Delta$. Afterwards, the functional form of the SGS model is inferred from the coarse-grained data using the Python implementation of SPIDER [24, 35]. SPIDER combines group representation theory (to construct term libraries which define the search spaces), weak formulation of differential equations (to reduce symbolic relations involving derivatives to a system of linear equations for a set of coefficients), and sparse regression (to infer equivariant functional relations corresponding to different irreducible representations of the relevant symmetry group, in this case, SO(2)) [25, 26, 27]. Finally, the relationship between the parameters of the effective field theory and the fundamental description (e.g., $\nu$) as well as the filter scale $\Delta$ are determined using a combination of dimensional analysis and scaling arguments.
@@
Results—To determine whether the SGS stress tensor can be represented by functions of the resolved fields $\bar{\bf u}$ and $\bar{p}$, we use them to construct the respective rank-0 and symmetric trace-free rank-2 term libraries and employ inhomogeneous regression. Inhomogeneous regression searches for sparse relations in the form $\tau=f(\bar{\bf u},\bar{p})$ with the known left-hand-side. Across all data sets, we consistently find the nonlinear gradient model (NGM2)
$\tau_{ij}$ $\approx c_{1}(\nabla_{k}\bar{u}_{i})(\nabla_{k}\bar{u}_{j})\equiv\tau_{ij}^{(2)}.$ (5)
Using dimensional analysis and the numerical values of the coefficients identified by SPIDER for different data sets, we find $c_{1}=\Delta^{2}/12$, consistent with the moment expansion of $L$. To obtain a more accurate approximation, we repeat the inhomogeneous regression for the difference $\tau_{ij}-\tau_{ij}^{(2)}$, which consistently yields the next order correction
$\tau_{ij}-\tau_{ij}^{(2)}$ $\approx c_{2}(\nabla_{k}\nabla_{m}\bar{u}_{i})(\nabla_{k}\nabla_{m}\bar{u}_{j})\equiv\tau_{ij}^{(4)}.$ (6)
Dimensional analysis and scaling yield $c_{2}=\Delta^{4}/288$, consistent with the moment expansion of $L+C-\tau^{(2)}$. Note that, for small $\Delta$, $\tau^{(4)}$ becomes small compared with $\tau^{(2)}$ and, as a result, is generally omitted by inhomogeneous regression for $\tau$. The iterative procedure described here reliably picks up this term. We will refer to the parameterization $\tau=\tau^{(2)}+\tau^{(4)}$ as the NGM4 model.
@@
Figure 2: The energy flux $\Pi$ describing the flow F2 shown in Figure 1(b) at $\nu=10^{-6}$ and cutoff scale $\Delta=\ell/64$ for DNS (a), the NGMR model (b), DS model (c), and DM model (d). All fields were calculated from DNS data.
As we show later, NGM4 is still not capable of accurately describing the energy flux. Hence, we repeat the iterative procedure to identify the next, even smaller correction $\tau^{(6)}=O(\Delta^{6})$. However, this time we find no robust parameterization for $\tau^{(6)}$ in terms of $\bar{\bf u}$ and $\bar{p}$ with even moderate accuracy, which means that $\tau^{(6)}$ must be expressed in terms of an additional tensor field describing the small scales. Since NGM4 already accounts for the contributions from $L$ and $C$ in the decomposition (4a), it is natural to assume that $\tau^{(6)}\approx R$ and treat $R$ as an additional variable in the effective field theory. This leads to the following parameterization of the SGS tensor
$\tau\approx\tau^{(2)}+\tau^{(4)}+R$ (7)
@@
which we refer to as the NGMR model. This model not only reproduces the individual components of the SGS stress tensor faithfully but also accurately captures the spatiotemporal structure of the energy flux $\Pi$ including backscatter (regions where $\Pi<0$) as illustrated in figure 2 for a particular flow and a particular choice of viscosity and the filter scale. For reference, the figure also shows the predictions of the widely used dynamic Smagorinsky (DS) model as well as the dynamic mixed (DM) model (see SI), which is considered to be one of the most accurate LES models [29].
In fact, the NGMR model succeeds on every metric and in every case we considered, while the phenomenological models rarely do. To measure the accuracy of an LES model in predicting a tensor variable, we employ a magnitude-aware correlation
between tensor fields $A$ (ground truth) and $B$ (prediction), where $\langle\cdot\rangle$ denotes spatial averaging, with $\mathcal{C}(A,B)=1$ precisely when $A=B$. In particular, $C_{\tau}=C(\tau_{\rm DNS},\tau_{\rm LES})$ and $C_{\Pi}=C(\Pi_{\rm DNS},\Pi_{\rm LES})$ describe the accuracy in predicting the components of the SGS tensor and the energy flux, respectively. The accuracy in predicting the net energy flux can be described in terms of the ratio $q_{\Pi}=\langle\Pi_{\text{LES}}\rangle/\langle\Pi_{\text{DNS}}\rangle$. All three quantities should be equal to unity (100%) if the prediction is perfect. As table 1 shows, both phenomenological models fail to predict the local fluxes for any flows and the net flux for flows F1 and F3. An unexpected observation is that a very accurate prediction of $\tau$ (as in the case of NGM4) does not guarantee accuracy in predicting the local and global energy fluxes. NGMR is the only model capable of predicting all these quantities across the board.
$\nu$ $10^{-4}$ $10^{-5}$ $10^{-6}$
ReRe $2\times 10^{4}$ $2\times 10^{4}$ $3\times 10^{5}$ $3\times 10^{5}$ $2\times 10^{5}$ $3\times 10^{6}$ $2\times 10^{6}$ $2\times 10^{6}$ $3\times 10^{7}$
model F1 F2 F3 F1 F2 F3 F1 F2 F3
DS 0.13% 0.01% 0.01% 0.24% 0.01% 0.01% 0.27% 0.01% 0.01%
DM 82.28% 99.71% 99.67% 73.52% 96.25% 99.67% 71.44% 96.17% 99.66%
$C_{\tau}$ NGM2 92.27% 99.27% 99.94% 89.21% 99.15% 99.93% 88.44% 99.14% 99.934%
NGM4 97.70% 99.93% 99.99% 95.94% 99.89% 99.99% 95.40% 99.89% 99.99%
NGMR 99.47% 99.98% 99.99% 99.34% 99.97% 99.99% 99.33% 99.97% 99.99%
DS 26.01% 24.86% 22.77% 26.28% 27.87% 24.05% 22.48% 30.63% 24.06%
DM 27.62% 18.00% 15.64% 29.63% 20.20% 13.41% 27.03% 21.08% 13.40%
$C_{\Pi}$ NGM2 0% 0% 0% 0% 0% 0% 0% 0% 0%
NGM4 83.49% 90.57% 99.44% 75.41% 83.81% 94.12% 72.26% 83.07% 93.98%
NGMR 96.85% 98.02% 99.85% 96.97% 97.52% 98.89% 97.12% 97.40% 98.83%
DS 131.76% 107.25% 31.53% 129.05% 106.27% 26.87% 128.74% 106.17% 26.86%
DM 134.12% 109.92% 35.03% 131.36% 108.93% 30.19% 131.04% 108.83% 30.18%
$q_{\Pi}$ NGM2 0% 0% 0% 0% 0% 0% 0% 0% 0%
NGM4 59.31% 69.83% 82.33% 49.13% 62.29% 58.60% 46.33% 60.78% 56.28%
NGMR 104.18% 104.63% 103.38% 103.81% 104.08% 104.94% 102.71% 103.93% 104.63%
Table 1: The accuracy, averaged in time over the dataset, of the SGS stresses and corresponding energy fluxes for flows F1–F3, with $R$ computed from DNS. To test the limits of the inferred parameterization, we purposely chose a coarse resolution with a cutoff scale $\Delta=\ell/64$ that is close to the size of the eddies in the flow F1.
@@
The inclusion of $R$ as a new field requires an additional governing equation in order to close the system. This equation can be inferred using rank-0 and symmetric trace-free rank-2 libraries constructed from $\bar{u}_{i}$, $\bar{p}$, and $R_{ij}$ and homogeneous regression, which searches for sparse relations in the form $f(R,\bar{\bf u},\bar{p})=0$. We robustly discover an evolution equation of the form
$\partial_{t}R_{ij}$ $\approx c_{1}\bar{u}_{k}\nabla_{k}R_{ij}+c_{2}(R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i})$
@@
Using a combination of dimensional analysis and numerical values of the coefficients for different data sets, we find $c_{1}=-1$, $c_{2}=1$ and $c_{3}=\nu$. For $c_{4}$, we find that it has both dimensions and numerical values consistent with the inverse of the eddy turnover time but exhibits no clear dependence on the parameters $\nu$ and $\Delta$. Therefore, we relax SPIDER’s assumption of translational invariance that requires $c_{4}$ to be independent of position and time and take it to be a function of the resolved flow. For instance, in the inviscid limit, rescaling the velocity of the initial condition by a constant factor changes the inverse time scale and, hence, $c_{4}$ by the same factor, so $c_{4}$ should be linear in $\bar{u}$. Because $c_{4}$ is a scalar, it should be expressible as a function of the energy $E=u_{i}u_{i}/2$ as well as the invariants of the tensor $\nabla_{i}\bar{u}_{j}$ [30]; in 2D, these are given by $|\bar{S}|=\sqrt{2\bar{S}_{ij}\bar{S}_{ij}}$ and $|\bar{\Omega}|=\sqrt{2\bar{\Omega}_{ij}\bar{\Omega}_{ij}}$, where ${\Omega}_{ij}$ is the antisymmetric component of $\nabla_{i}u_{j}$. Dimensional arguments would then require $c_{4}=\Delta^{-1}\sqrt{E}f(\Delta|\bar{\Omega}|/\sqrt{E},\Delta|\bar{S}|/\sqrt{E},\cdots)$, where $f$ is an arbitrary function. A general relation of this form can be inferred from the data using, e.g., symbolic regression. However, a good approximation can be found using a series expansion of $f$:
@@
$\partial_{t}R_{ij}$ $+\bar{u}_{k}\nabla_{k}R_{ij}\approx R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i}$ (11)
@@
Using $C_{R}=\mathcal{C}(\partial_{t}R_{\text{DNS}},\partial_{t}R_{\text{NGMR}})$ as a metric, we find equation (11) to accurately describe all flow regimes for sufficiently small $\Delta$ (cf. figure 3). As expected, the accuracy deteriorates as $\Delta$ approaches the scale of coherent structures, since the coarse-grained description is not meant to model these.
Figure 3: The accuracy of the evolution equation (11), quantified by the correlation $C_{R}$, as a function of $\Delta/\ell$ for F1 (blue), F2 (orange) and F3 (yellow) at $\nu=10^{-6}$.
Conclusions—Taken together, the closed system of equations (3), (7) and (11) defines an effective field theory which describes incompressible fluid turbulence up to the cutoff scale $\Delta$. By construction, this theory is explicitly equivariant with respect to translations and rotations but also turns out to be explicitly Galilean-invariant, as it should. Not only is the resulting description more accurate and general than state-of-the-art LES models—at least in two spatial dimensions—it is explicit and easily interpretable, which allows one to make specific predictions. For example, at scales $\Delta$ small compared to the size of large-scale coherent structures, the energy flux is dominated by the contribution from the $C$ tensor and scales as $\Delta^{4}$, which is not directly implied by the fundamental model (1). Our data-driven framework also identifies two methodological problems with traditional modeling approaches. Accurate description of inter-scale fluxes requires introduction of additional fields describing subgrid scales which are missing in all LES models. Moreover, these fields must have a tensor structure, unlike the most common RANS models such as $k-\epsilon$, $k-\omega$, and Spalart-Allmaras, which all employ scalar fields.
The traditional modeling approach has required decades of slow iterative progress and numerous—and often poorly justified—phenomenological assumptions to develop the current generation of LES models. In contrast, the framework presented here is both far more efficient and produces more accurate SGS models. Most importantly, it requires a minimal number of ingredients: the fundamental model (which could also be derived in a data-driven way [26]), some of the fundamental symmetries of the problem, dimensional analysis, and no specific phenomenological assumptions. Hence, it should readily generalize to other multiscale problems described by equations with quadratic nonlinearities, as discussed in the SI.